# Burgers equation from non-thermal stationary states in nearly-integrable gases

### v1

@@ III.3 Stochastic velocity swap models
where constant $A$ is fixed by normalization $\int{\rm d}\lambda\rho_{\mathrm{p}}^{\rm st}(\lambda)=\varrho$. The algorithm which implements classical model as well as Pauli principle in scatterings is presented in Appendix B. The models (37) and (38) are introduced as benchmarks for our effective, hydrodynamic theory, which can be efficiently simulated. In Fig. 2 we present numerical simulations of the classical version of the model, confirming the existence of athermal stationary states as well as the correctness of kinetic equation (37).
Figure 2: Relaxation dynamics of kinetic energy per particle $e(t)$ in classical stochastic velocity swap model. The molecular simulation agrees very well with kinetic equation (37). Inset: initial and final distributions of velocities. Solid lines are analytical distributions, in particular the final state agrees very well with (20). We consider homogenous system of $N=1600$ particles with $\varrho=1,\varrho^{(b)}=100$ and with $\gamma=0.001$. The initial system state is thermal with temperature $T=0.08$, whereas the bath system state is asymmetric Bragg state $\rho_{p}^{(b)}\propto e^{-(\lambda+\lambda_{B})^{2}/\sigma^{2}}+be^{-(\lambda-\lambda_{B})^{2}/\sigma^{2}}$ with $\sigma=0.2$, $\lambda_{B}=0.4$ and $b=0.5$. The results were averaged over 1000 realizations.
@@ IV Hydrodynamic equation
The effective hydrodynamic equation is obtained using the Chapman–Enskog method known from the kinetic theory [31]. Before going into details, let us start with summarizing the most important results of this section.
The equation for density which we find on large scales in the most general case takes the form (2). Expressions for the parameter $u$, related to the local center-of-mass velocity of the system, and the diffusion coefficient $\mathcal{D}(\varrho)$ are
@@
Here $a(\lambda,\varrho)$ is a solution to the Chapman-Enskog equation detailed below. The parameter $u(\varrho)$ vanishes for symmetric stationary states $\rho^{\rm st}$, leading to diffusion equation (1). For asymmetric state $\rho^{\rm st}$, which cannot be brought to symmetric states by Galilean boost, there are two cases. For classical statistics of system particles, $u(\varrho)=\text{const}$ and it can be removed through boost. For quantum statistics, $u(\varrho)$ is density-dependent, leading to Burgers equation.
We move now to the Chapman-Enskog method. In papers [42, 41] similar computation is performed for nearly-integrable system with total number of particles, momentum and energy conserved. At the foundations of the Chapman–Enskog method lies the separation of scales between the microscopic collision integral relaxation time $\tau_{\mathcal{I}}$ and macroscopic relaxation due to $\partial_{x}\left(v_{\rho_{\mathrm{p}}}\,\rho_{\mathrm{p}}\right)$ term in the GHD equation. The separation of scales gives us a suitable small parameter in which various quantities can be expanded. To accommodate the GHD diffusion in the (perturbative) Chapman–Enskog method we must additionally assume separation between characteristic spatial scale of the GHD diffusion and the characteristic spatial scale of $\rho_{\mathrm{p}}$ [42]. More precisely, we demand
@@
where $\ell_{\rm h}$ is macroscopic length characterizing spatial variations of $\rho_{\mathrm{p}}$ and $\langle v\rangle$ is the characteristic velocity of the state $\rho_{\mathrm{p}}$. By $\ell_{\mathrm{mic}}$ we denoted characteristic spatial scale of the GHD diffusion process. The introduced small parameters are called Chapman–Enskog parameters. In classical kinetic theory $\delta_{\mathcal{I}}$ is known as Knudsen number. For more discussion on the second Chapman–Enskog parameter we refer to [42]. Another important aspect of the Chapman–Enskog method is the assumption that time enters $\rho_{\mathrm{p}}$ only implicitly through the time-dependence of conserved charge, i.e.,
@@
IV.1 Evaluation of Chapman–Enskog procedure
@@
where in the second equality we used that, for a Galilean invariant system, $v_{\rho}$ satisfies integral equation (11), supplemented with a fact that $\mathcal{T}(\lambda)$ is symmetric [41]. In obtaining (43) we also used the Markovian property of diffusion matrix, discussed in Appendix.
To see how Chapman–Enskog parameters (41) appear in the GHD-Boltzmann equation (15) we rewrite the former in the dimensionless variables [42, 41, 31] $\bar{t}=t/\tau_{\mathrm{h}}$ and $\bar{x}=x/\ell_{\mathrm{h}}$ where $\tau_{\mathrm{h}}=\ell_{\mathrm{h}}/\langle v\rangle$ is the characteristic equilibrium relaxation time. In the new variables equation (15) takes the form
@@
where $\overline{v}_{\rho}=v_{\rho}/\langle v\rangle$, $\overline{\mathfrak{D}}_{\rho}=\mathfrak{D}_{\rho}/(\ell_{\mathrm{mic}}\,\langle v\rangle)$ and $\bar{\mathcal{I}}=\tau_{\mathcal{I}}\mathcal{I}$ are respectively velocity, diffusion operator and collision integral in the natural units and $\delta_{\mathfrak{D}}=l_{\rm mic}/l_{h}$ and $\delta_{Q}=\tau_{\mathcal{I}}\langle v\rangle/l_{h}$ are two small parameters. In what follows we drop $\bar{\cdot}$ symbols from the notation and keep $\delta$ parameters. After truncating to a given order, we will restore the original units. Since we assumed $\rho_{\mathrm{p}}$ to have the normal form (42) we can use the particle conservation law (43) to write the equation above without explicit time dependence:
@@
Solution of the Chapman-Enskog procedure will determine $G(x,\varrho)$ in terms of $\varrho$ thus closing the equation.
Next we proceed with expansion in small Chapman–Enskog parameters (41). According to the usual scheme, we keep the full dependence on $\delta_{\mathcal{I}}$ and $\delta_{\mathfrak{D}}$ in the conserved quantity $\varrho$ [31] and expand
@@
In order to solve (46) perturbatively in Chapman–Enskog parameters we also have to expand $\mathcal{I}[\rho_{\mathrm{p}}]$, $v_{\rho_{\mathrm{p}}}$ and $\mathfrak{D}_{\rho_{\mathrm{p}}}$ in terms of Chapman–Enskog parameters. Proceeding in analogy to (48) we denote expansion coefficients as $\mathcal{I}^{(i,j)}$, $v^{(i,j)}$ and $\mathfrak{D}^{(i,j)}$. Since they are all functionals of $\rho_{\mathrm{p}}$ we need to relate each of them with $\rho_{\mathrm{p}}^{(i,j)}$. This can be done by performing functional Taylor expansion of $\mathcal{I}^{(i,j)}$, $v^{(i,j)}$ and $\mathfrak{D}^{(i,j)}$ around $\rho_{\mathrm{p}}^{(0,0)}$, since $\rho_{\mathrm{p}}^{(0,0)}$ is the only factor of order unity in the expansion of $\rho_{\mathrm{p}}$.
Keeping full dependence on Chapman–Enskog parameters in $\varrho$, we are formally treating $\varrho$ as a quantity of order unity [31]. Since the only quantity of order unity in expansion of $\rho_{\mathrm{p}}$ is $\rho_{\mathrm{p}}^{(0,0)}$, for the theory to be consistent we require [31]
@@ IV.2 Euler scale
This condition fixes a state in each fluid cell to be stationary. From now on, we will denote the stationary state $\rho_{\mathrm{p}}^{(0,0)}$ as $\rho_{\mathrm{p}}^{\mathrm{st}}$. In the present context, we assume that the stationary states are parametrized by their density. Therefore, equation (50) has a family of solutions $\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho)$ where $\varrho=\int{\rm d}\lambda\rho_{\mathrm{p}}^{\mathrm{st}}$.
To find the Euler scale equations we recall that the density $\varrho$ varies in space and the quasi-particle distribution is given by $\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho(x,t))$. We can evaluate now the leading contribution to $G$ thus closing the continuity equation (47). We have
@@
Let us now gain some intuition on the velocity parameter $u$. First of all it can be shown that under the Galilean boost with velocity $w$, it transforms indeed as velocity: $u\rightarrow u+w$. This renders the Euler-scale equation Galilean invariant as it should be.
Furthermore, for symmetric states (this includes thermal states) $u=0$. This follows from the observations that the derivative can be taken in front of the integral and that the first moment of the thermal distribution is 0. A Galilean boost makes the first moment finite but density independent. Therefore, for standard collision integral with only thermal stationary term there is no dynamics at the Euler scale. A result known from the standard kinetic theory. However, for non-thermal stationary states $u$ can be non-zero. This requires the stationary state to be asymmetric in $\lambda$ and this asymmetry cannot be removed by a boost.
@@ IV.3 Diffusive scale
The second equation implies that $\rho^{(0,1)}$ is in the kernel of $\Gamma^{\mathrm{st}}$. This implies that $\rho^{(0,1)}=0$. We are left with the equation for $\rho^{(1,0)}$. We substitute for $G^{(0,0)}$ and use the chain rule in the last term to obtain
@@
By combining the Euler scale equation with the contribution from the diffusive scale, we finally obtain the Burgers (2). The diffusion coefficient $\mathcal{D}(\varrho)$ is invariant under the Galilean transformations, see Appendix C.3. Note that there is no contribution from the diffusion present in the GHD as $G^{(0,1)}=0$. This concludes the derivation of the hydrodynamic equation for density.
@@ V Applications
In the following section we apply the Chapman-Enskog formalism to coupled Bose gases, stochastic velocity swap models and to the so-called relaxation time approximation of collision operator.
@@ V.1 Relaxation time for two tubes system
We turn now to a more detailed picture of the collision integral which arises in a quasi-1d systems of two 1d gases that are coupled with a density-density interaction. In the density imbalance limit, as discussed in Section III, the collision integral $Q[\rho_{\rm p}]$ in the minority tube, assumed to be tube 1, is given by the Masters equation (24). The dynamic structure factor of the bath tube is given by [20]
@@
In the density imbalance limit evolution of the environmental tube is frozen i.e., $\partial_{t}\rho_{\mathrm{p}}^{(2)}=\mathcal{I}_{21}[\rho_{\mathrm{p}}^{(2)},\rho_{\mathrm{p}}^{(1)}]\approx 0$. Therefore we are free to choose $\rho_{\mathrm{p}}^{(2)}$ to be any stationary state. We will denote the state in the environmental tube as $\rho_{\mathrm{p}}^{(b)}$ and will refer to the particle density in the evolving tube $\rho_{\mathrm{p}}^{(1)}$ simply as $\rho_{\mathrm{p}}$. Having fixed the state in the bath tube, let’s assume that the state of the system tube is close to a stationary state
@@
Detailed computation of $\Gamma^{\mathrm{st}}$ is presented in the appendix D. Given that $\mathcal{I}_{12}$ has only one collision invariant, operator $\Gamma^{\mathrm{st}}$ has exactly one left null vector, corresponding to the conservation of the number of particles, see (16). Denoting eigenvalues as $\gamma_{i}$ we have $0=\gamma_{0}>\gamma_{1}\geq\gamma_{2}\geq\ldots$. Except for the non-decaying zeroth mode, each eigenmode of $\Gamma^{\mathrm{st}}$ will decay with the characteristic time $1/|\gamma_{i}|$. The characteristic long-term relaxation time can be identified with a gap between zero mode and smallest non-zero eigenvalue of $\Gamma^{\mathrm{st}}$ [21].
@@
Numerical estimation of the relaxation time is discussed in the Appendix E. This gives a conservative thermalisation scale and observables that do not couple to the slowest decaying charge will thermalise at faster rate.
@@ V.2 Relaxation Time Approximation
Given the complex nature of the collision integral, to be able to say more about the evolution without restricting to the specific system we need to make some approximation. One of the simplest approximation to apply is the Generalized Relaxation Time Approximation [21]. It amounts to reducing the structure of the operator $\delta\mathcal{I}/\delta\rho_{\mathrm{p}}$ to a scalar multiplication
@@
where multiplication by $\tau_{\mathrm{RTA}}^{-1}$ enforces thermalization at proper time scale. We only consider the case when the relaxation time is independent of $\varrho$. In the equation above $\rho_{\mathrm{p}}^{\mathrm{st}}$ is a stationary state towards which $\rho_{\mathrm{p}}$ evolves. If the only non-trivially conserved charge is the total number of particle then $\rho_{\mathrm{p}}^{\mathrm{st}}$ is related to $\rho_{\mathrm{p}}$ via $\int{\rm d}\lambda\,\rho_{\mathrm{p}}^{\mathrm{st}}=\int{\rm d}\lambda\,\rho_{\mathrm{p}}$. This relation makes $\rho_{\mathrm{p}}^{\mathrm{st}}[\rho_{\mathrm{p}}]$ highly non-linear functional of $\rho_{\mathrm{p}}$ [21]. Our ultimate goal in this section is to compute the diffusion coefficient (62) for the collision integral in the relaxation time approximation. First thing to do is to compute linearized collision integral (55):
@@
$j_{1}=\int{\rm d}\lambda\,h_{1}(\lambda)v(\lambda)\rho_{\mathrm{p}}^{\mathrm{st}}.$ (74)
@@
$\mathcal{D}_{\rm RTA}^{\rm th}=\frac{\tau_{\mathrm{RTA}}}{\varrho\kappa_{T}},\qquad\frac{1}{\kappa_{T}}=\varrho\left(\frac{\partial P}{\partial\varrho}\right)_{T}.$ (77)
The diffusion coefficient $\mathcal{D}_{\rm RTA}^{\rm th}$ is positive since all the quantities on the right hand side of (77) are positive. The positivity of $\kappa_{T}$ reflects the stability of thermal equilibrium. Here $P$ is the 11d pressure and it has the same dimensions as force.
In the upper part of Fig. 3 we present comparison of the diffusion coefficient computed in the RTA with the diffusion coefficient computed using the full collision integral for the two (Lieb–Liniger) tubes system  [^1] The inter-tube interaction potential was chosen to be a Gaussian potential $V(x)=V_{0}e^{-\left(x/x_{r}\right)^{2}}/(\sqrt{\pi}\,x_{r})$ with $x_{r}=2$. Constant $V_{0}$ can be expressed as $\tau_{\mathcal{I}}^{-1}=2A_{0}^{2}m/\hbar$ [52]. For the purpose of numerical computations we set $m=1$. Characteristic relaxation time is extracted from spectrum of the linearized collision integral in the limit $\varrho\to 0$.. For more detailed description of the numerical procedure we refer to the Appendix E. For the analyzed system we observe that RTA gives a good approximation of the diffusion coefficient. In the lower panel figure 3 we present numerically-computed $\nu$ for the reference tube in an asymmetric Bragg split state, i.e., state of the form
$\rho_{\mathrm{p}}^{(b)}(\lambda)=\eta\,\rho_{\mathrm{p}}^{\mathrm{th}}(\lambda-\Delta)+(1-\eta)\,\rho_{\mathrm{p}}^{\mathrm{th}}(\lambda+\Delta),$ (78)
where $\rho_{\mathrm{p}}^{\mathrm{th}}$ is a Gibbs thermal state, $\eta\in[0,1]$ is an imbalance parameter and $\Delta$ fixes the peak separation. Numerically computing $u^{\prime}$, we find that its value varies over several orders of magnitude, depending on the environmental tube’s state. Largest values of $u^{\prime}$ were found for small temperatures and small Bragg peaks overlap region.
@@
from which the diffusion coefficient can be read off. Comparing with the result of the Chapman-Enskog theory within the RTA we find the relation between the friction and the relaxation time given by $\tau_{\mathrm{RTA}}=\frac{m}{\zeta}$. This provides a way to estimate the relaxation time from the kinematics of a single particle.
Figure 3: In the upper plot, comparison of diffusion coefficient stemming from RTA with non-approximated diffusion for the two-tubes system in a thermal state ($T=4.0$) in Tonks–Girardeau limit. In the lower plot we show the function $\nu=u^{\prime}/\mathcal{D}_{\rm{RTA}}$ for an asymmetric Bragg split. Blue color represents a system in a small temperature ($T=1.0$), with the peak separation chosen in a way so that they overlap over a small region in $\lambda$. For the red-colored system we increased the temperature ($T=10.0$), simultaneously increasing peak separation. By $\tau_{\rm{RTA,i}}$ we understand the characteristic relaxation time of the system in the blue/orange state

### v2

@@ III.3 Stochastic velocity swap models
where constant $A$ is fixed by normalization $\int{\rm d}\lambda\rho_{\mathrm{p}}^{\rm st}(\lambda)=\varrho$. The algorithm which implements classical model as well as Pauli principle in scatterings is presented in Appendix C. The models (38) and (39) are introduced as benchmarks for our effective, hydrodynamic theory, which can be efficiently simulated. In Fig. 2 we present numerical simulations of classical version of the model in homogenous setting, confirming the existence of athermal stationary states as well as the correctness of kinetic equation (38).
Figure 2: Relaxation dynamics of kinetic energy per particle $e(t)$ in classical stochastic velocity swap model. The molecular simulation agrees very well with solution of kinetic equation (38). Inset: initial and final distributions of velocities. Solid lines are analytical distributions, in particular the microscopic final state agrees very well with analytical prediction (20). We consider homogenous system of $N=1600$ particles with $\varrho=1,\varrho^{(b)}=100$ and with $\gamma=0.001$. The initial system state is thermal with temperature $T=0.08$, whereas the bath system state is asymmetric Bragg-split state given by $\rho_{\mathrm{p}}^{(b)}\propto e^{-(\lambda+\lambda_{B})^{2}/\sigma^{2}}+be^{-(\lambda-\lambda_{B})^{2}/\sigma^{2}}$ with $\sigma=0.2$, $\lambda_{B}=0.4$ and $b=0.5$. The results were averaged over 1000 realizations.
@@ IV Hydrodynamic equation
The effective hydrodynamic equation is obtained using the ChE method known from the kinetic theory [49]. Before going into the details of the method, let us start with summarizing the most important results of this section.
The equation for density $\varrho(x,t)$ which we find on large scales in the most general case takes the form (2). Expressions for the advective coefficient $u(\varrho)$, related to the local center-of-mass velocity of the system, and the diffusion coefficient $\mathcal{D}(\varrho)$ are
@@
Here $a(\lambda,\varrho)$ is a solution to the ChE equation detailed below. The parameter $u(\varrho)$ vanishes for parity-symmetric stationary states $\rho_{\mathrm{p}}^{\rm st}$, leading to diffusion equation (1). For parity-asymmetric state $\rho_{\mathrm{p}}^{\rm st}$, which cannot be brought to symmetric states by Galilean boost, there are two cases. For classical statistics of system particles, $u(\varrho)=\text{const}$ and it can be removed through boost. For quantum statistics, $u(\varrho)$ is density-dependent, leading to Burgers equation.
We move now to the ChE method. In papers [61, 60] similar computation is performed for nearly-integrable system with total number of particles, momentum and energy conserved. At the foundations of the ChE method lies the separation of scales between the microscopic collision integral relaxation time $\tau_{\mathcal{I}}$ and macroscopic relaxation due to $\partial_{x}\left(v_{\rho_{\mathrm{p}}}\,\rho_{\mathrm{p}}\right),\partial_{x}(\mathfrak{D}_{\rho}\partial_{x}\rho_{\mathrm{p}})$ terms in the GHD equation. The separation of scales gives us a suitable small parameter in which various quantities can be expanded. To accommodate the GHD diffusion in the (perturbative) ChE method we must additionally assume separation between characteristic spatial scale of the GHD diffusion and the characteristic spatial scale of $\rho_{\mathrm{p}}$ [61]. More precisely, we demand
@@
where $\ell_{\rm h}$ is hydrodynamic length characterizing spatial variations of $\rho_{\mathrm{p}}$ and $\langle v\rangle$ is the characteristic velocity of the state $\rho_{\mathrm{p}}$. By $\ell_{\mathrm{mic}}$ we denoted characteristic spatial scale of the GHD diffusion process. The introduced small parameters $\delta_{\mathcal{I}},\delta_{\mathfrak{D}}$ are called ChE parameters. In classical kinetic theory $\delta_{\mathcal{I}}$ is known as Knudsen number. For more discussion on the second ChE parameter we refer to [61]. Another important aspect of the ChE method is the assumption, that time enters $\rho_{\mathrm{p}}$ only implicitly through the time-dependence of conserved charge, i.e.,
@@
IV.1 Evaluation of ChE procedure
@@
where in the second equality we used that, for a Galilean invariant system, $v_{\rho}$ satisfies integral equation (11), supplemented with a fact that $\mathcal{T}(\lambda)$ is symmetric [60]. In obtaining (44) we also used the Markovian property of diffusion matrix, discussed in Appendix D.1. Obviously, at this point (44) is not closed as the current depends on the whole distribution $\rho_{\mathrm{p}}$. To make some progress we have to take advantage of smallness of ChE parameters.
To see how ChE parameters (42) appear in the GHD-Boltzmann equation (15) we introduce dimensionless variables [61, 60, 49] $\bar{t}=t/\tau_{\mathrm{h}}$ and $\bar{x}=x/\ell_{\mathrm{h}}$ where $\tau_{\mathrm{h}}=\ell_{\mathrm{h}}/\langle v\rangle$ is the characteristic hydrodynamic time scale. In the new variables equation (15) takes the form
@@
where $\overline{v}_{\rho}=v_{\rho}/\langle v\rangle$, $\overline{\mathfrak{D}}_{\rho}=\mathfrak{D}_{\rho}/(\ell_{\mathrm{mic}}\,\langle v\rangle)$ and $\bar{\mathcal{I}}=\tau_{\mathcal{I}}\mathcal{I}$ are respectively velocity, diffusion operator and collision integral in the natural units. In what follows we drop $\bar{\cdot}$ symbols from the notation and keep $\delta$ parameters. After truncating to a given order, we will restore the original units. Since we assumed $\rho_{\mathrm{p}}$ to have the normal form (43) we can use the particle conservation law (44) to write the equation above without explicit time dependence:
@@
Solution of the ChE procedure will determine $G(x,\varrho)$ in terms of $\varrho$ thus closing the equation.
Next we proceed with expansion in small ChE parameters (42). According to the usual scheme, we expand the density $\rho_{\mathrm{p}}$ and $G$ in the two ChE parameters  [49] and expand
@@
In order to solve (47) perturbatively in ChE parameters we also have to expand $\mathcal{I}[\rho_{\mathrm{p}}]$, $v_{\rho_{\mathrm{p}}}$ and $\mathfrak{D}_{\rho_{\mathrm{p}}}$ in terms of ChE parameters. Proceeding in analogy to (49) we denote expansion coefficients as $\mathcal{I}^{(i,j)}$, $v^{(i,j)}$ and $\mathfrak{D}^{(i,j)}$. Since they are all functionals of $\rho_{\mathrm{p}}$ we need to relate each of them with $\rho_{\mathrm{p}}^{(i,j)}$. This can be done by performing functional Taylor expansion of $\mathcal{I}^{(i,j)}$, $v^{(i,j)}$ and $\mathfrak{D}^{(i,j)}$ around $\rho_{\mathrm{p}}^{(0,0)}$, since $\rho_{\mathrm{p}}^{(0,0)}$ is the only factor of order unity in the expansion of $\rho_{\mathrm{p}}$.
Finally, and contrary to other quantities, we do not expand the conserved density $\varrho$. Since the only quantity of order unity in expansion of $\rho_{\mathrm{p}}$ is $\rho_{\mathrm{p}}^{(0,0)}$, for the theory to be consistent we require [49]
@@ IV.2 Euler scale
This condition fixes a state in each fluid cell to be stationary state of collision operator. From now on, we will denote the stationary state $\rho_{\mathrm{p}}^{(0,0)}$ as $\rho_{\mathrm{p}}^{\mathrm{st}}$. In the present context, we assume that the stationary states are parametrized by their density. Therefore, equation (51) has a family of solutions $\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho)$ where $\varrho=\int{\rm d}\lambda\rho_{\mathrm{p}}^{\mathrm{st}}$.
To find the Euler scale equations we recall that the density $\varrho$ varies in space and the quasiparticle distribution is given by $\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda,\varrho(x,t))$. We can evaluate now the leading contribution to $G$ thus closing the continuity equation (48). We have
@@
Let us now gain some intuition on the advective coefficient $u$. First of all, it can be shown that under the Galilean boost with velocity $w$, it transforms indeed as velocity: $u\rightarrow u+w$ (see Appendix D.3). This renders the Euler-scale equation Galilean invariant as it should be.
Furthermore, for parity-symmetric states (this includes thermal states) $u=0$. This follows from the observations that the derivative can be taken in front of the integral and that the first moment of a symmetric distribution is 0. A Galilean boost makes the first moment finite but density independent. Therefore, for standard collision integral with only thermal stationary term there is no dynamics at the Euler scale. A result known from the standard kinetic theory. However, for non-thermal stationary states $u$ can be non-zero. This requires the stationary state to be asymmetric in $\lambda$ and this asymmetry cannot be removable by a boost.
@@ IV.3 Diffusive scale
The second equation implies that $\rho^{(0,1)}$ is in the kernel of $\Gamma^{\mathrm{st}}$. This implies that $\rho^{(0,1)}=0$, due to constitutive relations. We are left with the equation for $\rho^{(1,0)}$. We substitute for $G^{(0,0)}$ and use the chain rule in the last term to obtain
@@
By combining the Euler scale equation with the contribution from the diffusive scale, we finally obtain the Burgers (2). The diffusion coefficient $\mathcal{D}(\varrho)$ is invariant under the Galilean transformations, see Appendix D.3. Note that there is no contribution from the diffusion present in the GHD as $G^{(0,1)}=0$. We also mention that the diffusion coefficient in general scales as an inverse of characteristic timescale of $\mathcal{I}$ meaning that it is parametrically big in the strength of integrability breaking. This concludes the derivation of the hydrodynamic equation for density.
@@ V Applications
In the following section we apply the ChE formalism to coupled Bose gases, stochastic velocity swap models and to the so-called relaxation time approximation of collision operator.
@@ V.1 Relaxation time for two tubes system
We turn now to a more detailed picture of the collision integral which arises in a quasi-1D systems of two 1D gases that are coupled with a density-density interaction. In the density imbalance limit, as discussed in Section III, the collision integral $\mathcal{I}[\rho_{\rm p}]$ in the minority tube, assumed to be tube 1, is given by the Master equation (24). The dynamic structure factor of the bath tube is given by [38]
@@
In the density imbalance limit evolution of the environmental tube is frozen i.e., $\partial_{t}\rho_{\mathrm{p}}^{(2)}=\mathcal{I}_{21}[\rho_{\mathrm{p}}^{(2)},\rho_{\mathrm{p}}^{(1)}]\approx 0$. Therefore we are free to choose $\rho_{\mathrm{p}}^{(2)}$ to be any stationary state. We will denote the state in the environmental tube as $\rho_{\mathrm{p}}^{(b)}$ and will refer to the particle density in the evolving tube $\rho_{\mathrm{p}}^{(1)}$ simply as $\rho_{\mathrm{p}}$. Let us assume that (homogenous) state of the system tube is close to a stationary state
@@
Detailed computation of $\Gamma^{\mathrm{st}}$ is presented in the appendix E. Given that $\mathcal{I}_{12}$ has only one collision invariant, operator $\Gamma^{\mathrm{st}}$ has exactly one left null vector, corresponding to the conservation of the number of particles, see (16). Denoting eigenvalues as $\gamma_{i}$ we have $0=\gamma_{0}>\gamma_{1}\geq\gamma_{2}\geq\ldots$. Except for the non-decaying zeroth mode, each eigenmode of $\Gamma^{\mathrm{st}}$ will decay with the characteristic time $1/|\gamma_{i}|$. The characteristic relaxation time can be identified with a gap between zero mode and smallest non-zero eigenvalue of $\Gamma^{\mathrm{st}}$ [39].
@@
Numerical estimation of the relaxation time is discussed in the Appendix F. This gives a conservative thermalisation scale and observables that do not couple to the slowest decaying charge will thermalise at faster rate.
@@ V.2 Relaxation Time Approximation
Given the complex nature of the collision integral, to be able to say more about the evolution without restricting to the specific system we need to make some approximation. One of the simplest approximation to apply is the RTA [39]. It amounts to the following approximation of collision integral
@@
where multiplication by $\tau_{\mathrm{RTA}}^{-1}$ enforces relaxation at proper time scale. We only consider the case when the relaxation time is independent of $\varrho$. In the equation above $\rho_{\mathrm{p}}^{\mathrm{st}}$ is a stationary state towards which $\rho_{\mathrm{p}}$ evolves. If the only non-trivially conserved charge is the total number of particle then $\rho_{\mathrm{p}}^{\mathrm{st}}$ is related to $\rho_{\mathrm{p}}$ via $\int{\rm d}\lambda\,\rho_{\mathrm{p}}^{\mathrm{st}}=\int{\rm d}\lambda\,\rho_{\mathrm{p}}$. This relation makes $\rho_{\mathrm{p}}^{\mathrm{st}}[\rho_{\mathrm{p}}]$ highly non-linear functional of $\rho_{\mathrm{p}}$ [39]. Our ultimate goal in this section is to compute the diffusion coefficient (63) for the collision integral in the relaxation time approximation. First thing to do is to compute linearized collision integral (56):
@@
$j_{1}=\int{\rm d}\lambda\,h_{1}(\lambda)v(\lambda)\rho_{\mathrm{p}}^{\mathrm{st}}(\lambda).$ (75)
@@
$\mathcal{D}_{\rm RTA}^{\rm th}=\frac{\tau_{\mathrm{RTA}}}{\varrho\kappa_{T}},\qquad\frac{1}{\kappa_{T}}=\varrho\left(\frac{\partial P}{\partial\varrho}\right)_{T},$ (78)
where $P$ is the 1D pressure. The diffusion coefficient $\mathcal{D}_{\rm RTA}^{\rm th}$ is positive since $\kappa_{T}>0$ reflecting the stability of thermal equilibrium.
In the upper part of Fig. 3 we present comparison of the diffusion coefficient computed in the RTA with the diffusion coefficient computed using the full collision integral for the two (Lieb–Liniger) tubes system  [^1] The inter-tube interaction potential was chosen to be a Gaussian potential $V(x)=V_{0}e^{-\left(x/x_{r}\right)^{2}}/(\sqrt{\pi}\,x_{r})$ with $x_{r}=2$. Constant $V_{0}$ can be expressed as $\tau_{\mathcal{I}}^{-1}=2A_{0}^{2}m/\hbar$ [68]. For the purpose of numerical computations we set $m=1$. Characteristic relaxation time is extracted from spectrum of the linearized collision integral in the limit $\varrho\to 0$.. For more detailed description of the numerical procedure we refer to the Appendix F. For the analyzed system we observe that RTA gives a good approximation of the diffusion coefficient. In the lower panel figure 3 we present numerically-computed $\nu$ for the bath tube in an asymmetric Bragg-split state, i.e., state of the form
$\rho_{\mathrm{p}}^{(b)}(\lambda)\propto\rho_{\mathrm{p}}^{\mathrm{th}}(\lambda+\lambda_{B})+b\rho_{\mathrm{p}}^{\mathrm{th}}(\lambda-\lambda_{B}),$ (79)
where $\rho_{\mathrm{p}}^{\mathrm{th}}$ is a Gibbs thermal state, $b\in[0,1]$ is an asymmetry parameter and $\lambda_{B}$ fixes the peak separation. Numerically computing $u^{\prime}$, we find that its value varies over several orders of magnitude, depending on the environmental tube’s state. We find that $u^{\prime}$ attains the largest values in the regime of low temperatures, with $\lambda_{b}$ being slightly smaller than $({\rm width\;of\;}\rho_{\mathrm{p}}^{\mathrm{th}})/2$, meaning that left and right peaks minimally overlap with each other.
@@
from which the diffusion coefficient can be read off. Comparing with the result of the ChE theory within the RTA we find the relation between the friction and the relaxation time given by $\tau_{\mathrm{RTA}}=\frac{m}{\zeta}$. This provides a way to estimate the relaxation time from the kinematics of a single particle.
Figure 3: Transport coefficients in the coupled LL models in the Tonks-Girardeau limit. Panel (a): comparison of diffusion coefficient stemming from RTA with non-approximated diffusion for the two-tubes system in a thermal state ($T=4.0$). In the panels (b) and (c) we show the Burgers parameter $\nu=\mathcal{D}_{\rm{RTA}}/u^{\prime}$ for a system in asymmetric Bragg-split state (79). In (b) we present a system in a small temperature ($T=1.0$), with asymmetry $b=9$ with the peak separation $\lambda_{B}=3.9$ chosen in a way so that they overlap over a small region in $\lambda$. For (c), we increased the temperature ($T=10.0$), simultaneously increasing peak separation to $\lambda_{B}=8.0$ with the same $b=9$.